pick only one of the method of overtime from the main portion of the assignment. randomly pick a regular rate of pay and hours work for the week, the hours should be greater than 40 hours, illustrate a method with all the calculation for all steps.
Answer by fractalier(6550)

A union plumber normally earns $60 an hour. His normal work week is 40 hours.
He earns 1 1/2 times his regular pay for overtime each week.
One week he worked 48 hours. How much dd he earn?
Well he earned 40 hours times $60 an hour for his regular pay and 8 hours times 1 1/2 times $60 and hour for his over time, or
Total = 40(60) + 8(1.5)(60) = 2400 + 720 = $3120
Wish I was a union plumber!